The rate law for the decomposition of N2O5 is

rate = k[N2O5]

If k = 1.0 x 10-5 s-1, what is the reaction rate when the N2O5 concentration is 0.0091 mol L-1?

2.The decomposition of acetaldehyde, CH3CHO, was determined to be a second order reaction with a rate constant of 0.0771 M-1 s-1. If the initial concentration of acetaldehyde is 0.301 M , what will the concentration be after selected reaction times?

a. What will the CH3CHO concentration be after 30.0 minutes?

b. What will the CH3CHO concentration be after 180 minutes?
Answer #1

1) Rate law: rate *[N2O,] we have k=1.0×10-5 s-1 [Ν,0,1- Thus, 0.0091 mol L-1 rate =小1,0.] = (1.0×10-5 s-1)(0.0091 moll = 9.1